# reference.bib

### v2

@@ Abstract
Stabilizing plasma dynamics is an important task in green energy generation via nuclear fusion. One common strategy is to introduce an external field to prevent the plasma distribution from developing turbulence. However, finding such external fields efficiently remains an open question, even for simplified models such as the Vlasov-Poisson (VP) system. In this work, we leverage two different approaches to build such fields: for the first approach, we use an analytical derivation of the dispersion relation of the VP system to find a range of reasonable fields that can potentially suppress instability, providing a qualitative suggestion. For the second approach, we leverage PDE-constrained optimization to obtain a locally optimal field using different loss functions. As the stability of the system can be characterized in several different ways, the objective functions need to be tailored accordingly. We show, through extensive numerical tests, that objective functions such as the relative entropy (KL divergence) and the $L^{2}$ norm result in a highly non-convex problem, rendering the global minimum difficult to find. However, we show that using the electric energy of the system as a loss function is advantageous, as it has a large convex basin close to the global minimum. Unfortunately, outside the basin, the electric energy landscape consists of unphysical flat local minima, thus rendering a good initial guess key for the overall convergence of the optimization problem, particularly for solvers with adaptive steps.
@@ 1 Introduction
Plasma is an ionized gas composed of free electrons, ions, and neutral particles [burm2012plasma, fitzpatrick2022plasma, frank2012plasma]. One application for plasma is to generate green energy from fusion [ichimaru1993nuclear, miyamoto1980plasma].
However, in fusion devices, it is well known that plasma gives rise to many instabilities: a small perturbation to the equilibrium states, due to internal or external influences or imperfections, can grow and disrupt the system, rendering fusion devices impractical. As a consequence, to fully utilize plasma’s capabilities, it is essential to study and manage these instabilities. There are many kinds of equilibrium states for plasma: we study the class of equilibrium states that are homogeneous in space, e.g., [van2021beam, zohm2015magnetohydrodynamic].
One strategy to control instability is to impose external electric or magnetic fields to force plasma to move along desired trajectories. To find this control, mathematically one can formulate the problem as a PDE-constrained optimization: conditioned on plasma PDE being satisfied, one looks for the control (external fields) that suppresses instability:
$\min_{H}$ $\quad\quad\quad\quad\quad\quad\quad\mathcal{J}(f[H])$ (1.1)
@@
s.t. $\,\left\{\begin{array}[]{lc}\partial_{t}f+v\cdot\nabla_{x}f-(E_{f}+H)\cdot\nabla_{v}f=0\,,\\ E_{f}=\nabla_{x}V_{f}\,,\\ \Delta V_{f}=1-\rho_{f}=1-\int f\,\mathrm{d}v\,,\end{array}\right.$
@@
where the objective function $\mathcal{J}$ is a functional that quantifies the instability in the plasma distribution $f$, and the PDE is the Vlasov-Poisson (VP) system, the most fundamental kinetic model that describes the dynamics of plasma. In the equation, $f(t,x,v)$ stands for the density of plasma particles on the phase space $(x,v)\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ at time $t\in\mathbb{R}_{+}$, $E_{f}(t,x)$ is the self-generated electric field that comes from the gradient of the self-generated potential $V_{f}(t,x)$, which is determined by the charge density $\rho_{f}(t,x)$ such that
$\rho_{f}(t,x)=\int f(t,x,v)\,\mathrm{d}v\,,$
The quantities, $\rho_{f}$, $V_{f}$ and $E_{f}$ are macroscopic quantities and are functions of space only. $H(x)$ is the control (external electric field) and the parameter to be tuned. The formulation looks for the optimal control (an external electric field) $H$ so that $\mathcal{J}$, the instability is minimized.
This formulation is straightforward and has been adopted in multiple studies, both theoretically [Einkemmer2025, Glass2003, Glass2012, Knopf2018, Knopf2019] and numerically [Einkemmer2024, albi2025instantaneous, kawata2019dynamic, Bartsch2024], with variations replacing the Vlasov-Poisson equation with the fluid counterpart magnetohydrodynamic (MHD) system [Bialek2001, Strait2014]. These works laid the foundations for understanding instability control in plasma dynamics, providing key insights into the mathematical structure and computational strategies required for effective stabilization.
However, most of the currently available work assume the instability is presented as the simple difference between the computed solution $f_{T}:=f(T,\cdot,\cdot)$ and the desired equilibrium state. Denote $f_{\text{eq}}$ the desired equilibrium state, in [Einkemmer2024], the authors studied the optimization solver for
$\mathcal{J}(f)=\frac{1}{2}\|f_{T}-f_{\text{eq}}\|_{L^{2}(x,v)}^{2}\,.$
Namely, the PDE equipped with the external field $H$ is ran up to time $t=T$, and one wishes to minimize the difference between the PDE generated solution $f_{T}$ and the desired state.
However, it is unclear this is a good form to characterize the instability. To suppress instability, ultimately, one hopes to minimize the self-generated electric field, so to prevent plasma from mixing. This means the instability should be characterized as a quantity in spatial domain only. The $L^{2}$ distance between PDE output and the desired distribution examines the stability on $(x,v)$ domain, which may be unnecessarily excessive. Different objective functions, formulated to quantify different characterizations of instability, promote different properties of the PDE solution, and achieve different suppressing mechanism. Indeed, in [Einkemmer2024], the authors plotted a scan over two parameters for the external field $H$, and observed a non-convex, even highly oscillatory pattern for the objective function. This indicates that one should be cautious in selecting the objective function that truly represents the instability to suppress.
Upon a good choice of the objective, one then needs to run optimization algorithms to find optimal control and solve (1.1). There are many off-the-shelf optimization solvers, each bringing about different features. Different solvers are compatible to different type of objectives. For instance, if the solvers are “local” in nature (such as gradient descent with fixed time-stepping), they would not be able to reconstruct optimal solution to (1.1) when the landscape is rugged. To overcome this difficulty, a mechanism needs to be built in the solver that allows samples to move out of local minima basin. One simple fix is to deploy line search. This allows a sample to explore the best parameter configuration in the gradient direction. Another possibility is to employ global solvers such as particle swarm-based method or genetic algorithms can be deployed too, as was done in [Einkemmer2024]. However, these global solvers usually require multiple parallel sample paths, and can be much more computationally demanding.
These observations compel us to numerically examine the following question
@@
What is a good metric to quantify and optimize for suppressing instabilities in a Vlasov-Poisson system?
More specifically, we are interested in finding alternative metrics to the brute-force $L^{2}$ as defined in [Einkemmer2024], that can potentially better quantify the instability, and give more convex objective function landscapes. Moreover, we are also interested in finding optimization solvers that are compatible with these metrics.
Different mathematical definition of “instability” (objective functions) show drastically different landscape for a given shared physics problem is not a new phenomenon, and is especially pronounced for wave-type problems. It is well-known that VP system, upon a small perturbation from an equilibrium state, undergoes turbulence or filementation, making the solution highly oscillatory in the phase space. Consequently, a large integer period of a shift in phase space can produce a much smaller mismatch compared to a small non-integer shift. In wave-type inverse problems, this phenomenon is termed cycle-skipping, and has garnered a great amount of research interests [Engquist2022, Symes1991, ChenDing2023], with many of the works proposing alternative metrics for designing objective functions. Similar to these efforts for wave-inversion, we conjecture a different objective function, potentially defined on a macroscopic quantity (such as $\rho_{f}$ or $E_{f}$), can give a better landscape.
This is indeed the case. Our numerical finding suggest that defining $\mathcal{J}$ through the distribution function $f$, either using the standard $L^{2}$ distance or relative entropy against a desired equilibrium state, leads to a non-convex objective function with a very rugged landscape. Conversely, setting $\mathcal{J}$ to be the self-generated electric energy, a macroscopic quantity, relaxes the fine-grained requirements imposed on $f$, and give a much convexified landscape. As a result, a solver that can see beyond local gradient is necessary if $\mathcal{J}$ is defined through $f$, but a simple gradient descent with fixed (small) time-stepping can quickly suppresses instability if $\mathcal{J}$ only sees macroscopic quantity.
To initiate the optimization procedure, we resort to PDE analysis for a good range of parameter configuration. More specifically, using Laplace and Fourier transforms, one can deduce the dispersion relation for the VP system in the linearized regime. The root analysis for the dispersion relation reveals the information on the frequency of the modes that blow up the fastest. In return, these analysis provide a good guess on the control $H$: The field should be designed to focus on suppressing these blowing-up modes. The analysis is only valid in the linear regime, but nevertheless pinpoint the range for finding parameters for the optimal field, and hence serves as the initial guess for running optimization.
@@
The paper is organized as follows. In Section 2 we properly define the PDE-constrained optimization problem that we aim to solve, along the different possible objective functions that we will test. We also present the forward solver that we will use to solve the forward problem and the two canonical examples that we will present in this work. In Section 3 we briefly present the dispersion relation and a linear stability analysis which gives us important information on the frequency and scaling that our control $H$ should have. In Section 4 we study the different landscapes that the different objectives have. In Section 5 we make substantial numerical simulations with different objectives, methods, and initializations of the parameters.
@@ 2 Preliminaries
In practice, plasma for fusion applications is confined within a circular domain. From a mathematical perspective, this corresponds to treating the $y,z$-domain as constant and expressing the 3D problem as a pseudo-1D problem with periodic boundary conditions, where $(x,v)\in[0,L_{x}]\times[-L_{v},L_{v}]$ with periodic boundary conditions in space and homogeneous Dirichlet boundary conditions in the velocity domain. Under this framework, the Vlasov-Poisson system can be simplified, resulting in the following formulation:
$\left\{\begin{array}[]{lc}\partial_{t}f+v\partial_{x}f-(E_{f}+H)\cdot\partial_{v}f=0\,,\\ E_{f}=\partial_{x}V_{f}\,,\\ -\partial_{xx}V_{f}=\int f_{\text{eq}}-f\,\mathrm{d}v:=\rho_{\text{p}}\,.\end{array}\right.$ (2.1)
where we decomposed $f=f_{\text{p}}+f_{\text{eq}}$ such that $f_{\text{p}}$ represents the component of the dynamics that has been perturbed and our perturbed charge density $\rho_{\text{p}}$ is defined as
$\rho_{\text{p}}(t,v):=\int f_{\text{eq}}(v)-f(t,x,v)\,\mathrm{d}v\,.$
@@
It is well known that equilibrium states which do not satisfy the Penrose condition [penrose1960electrostatic] are classified as unstable. In such cases, even a small perturbation to the equilibrium state can lead to exponential growth of the self-generated electric field $E_{f}$. In this work, we study two canonical examples of such unstable equilibria that violate the Penrose condition: the Two-Stream equilibrium (Section 2.3.1) and the Bump-on-Tail equilibrium (Section 2.3.2).
@@ 2.1 PDE-constrained optimization problem
$\min_{H}$ $\quad\quad\quad\mathcal{J}(f[H])$ (2.3)
@@
where our objective function will be either of the following three
$\mathcal{J}(f[H])=\frac{1}{2}\|f_{T}[H]-f_{\text{eq}}\|^{2}_{L^{2}(x,v)}=\frac{1}{2}\int_{-L_{v}}^{L_{v}}\int_{0}^{L_{x}}(f[H](T,x,v)-f_{\text{eq}}(v))^{2}\,\mathrm{d}x\,\mathrm{d}v\,,$ ($L^{2}$)
@@
$\mathcal{J}(f[H])=\int_{0}^{T}\mathcal{E}_{f[H]}(t)\mathrm{d}t=\int_{0}^{T}\int_{0}^{L_{x}}[E_{f[H]}(t,x)]^{2}\,\mathrm{d}x\,\mathrm{dt}\,,$ (EE)
where $E_{f[H]}$ and $\mathcal{E}_{f[H]}$ are as defined in (2.1) and (2.2) respectively.

### v3

# What metric to optimize for suppressing instability in a Vlasov-Poisson system?

@@ Abstract
Stabilizing plasma dynamics is an important task in green energy generation via fusion. A common strategy is to introduce an external field to prevent the plasma distribution from becoming turbulent. However, finding such external fields efficiently remains an open question, even for simplified models such as the Vlasov-Poisson (VP) system. In this work, we present an integrated method where we first perform an analytical derivation of the VP system’s dispersion relation to construct a high-quality initial guess for the stabilizing field. This analytically-derived field is then used to initialize a PDE-constrained optimization loop that refines the control to a local optimum. Through extensive numerical experiments, we demonstrate that objective functions evaluated only at the target time—when stable plasma is desired—lead to highly non-convex optimization landscapes, making the global minimum difficult to locate. This behavior arises regardless of the choice of loss function (e.g., KL divergence or electric energy). In contrast, integrating the loss function over time yields a landscape with a convex basin near the global minimum, facilitating convergence. Furthermore, when using electric energy as the objective, the landscape outside this basin is dominated by flat, unphysical local minima, highlighting the critical importance of our analytical approach for generating an initial guess that lies within the global minimum’s basin of attraction.
@@ 1 Introduction
Plasma, an ionized gas composed of electrons, ions, and neutral particles, is often referred to as the fourth state of matter [10, 23, 24]. One of its most compelling applications is the generation of clean energy through nuclear fusion [30, 37]. A principal obstacle to realizing practical fusion energy is the inherent instability of plasma dynamics: in many configurations, small perturbations to an equilibrium state can grow exponentially, disrupting the system and rendering a fusion device inoperable. Consequently, developing robust strategies to suppress these instabilities is a critical area of research.
These instabilities are typically mitigated using control strategies, wherein an external actuator—such as a magnetic or electric field—is applied to the plasma to counteract the growth of disruptive modes. Control strategies can be broadly classified by their temporal nature. Dynamic control involves real-time feedback, where the actuator responds continuously to the evolving state of the plasma. This paradigm has been the subject of extensive research, employing both traditional optimal control techniques [1, 2, 7, 18] and modern reinforcement learning (RL) methods [16, 46]. However, dynamic approaches face a fundamental limitation: a timescale mismatch between the rapid growth of certain instabilities and the slower response time of physical actuators [38]. In response to this challenge, static control has recently regained significant interest from the community. This paradigm involves designing a time-independent control to dampen a small initial perturbation, with the goal of stabilizing the system at an ulterior time $T$ [19, 35, 36]. This work focuses on the design of such static controls.
A common approach for designing a static control is to impose an external electric field, $H(x)$, to guide the plasma towards a desired stable distribution [19]. This task can be formulated as a PDE-constrained optimization problem:
$\min_{H}$ $\quad\mathcal{J}(f[H])$ (1.1)
@@
s.t. $\quad\left\{\begin{array}[]{l}\partial_{t}f+v\cdot\nabla_{x}f-(E_{f}+H)\cdot\nabla_{v}f=0\,,\\ E_{f}=\nabla_{x}V_{f}\,,\\ \Delta_{x}V_{f}=1-\rho_{f}=1-\int f\,\mathrm{d}v\,.\end{array}\right$.
@@
Here, the constraint is the Vlasov-Poisson (VP) system, a fundamental kinetic model for collisionless plasma. The function $f:(t,x,v)\in\mathbb{R}_{+}\times\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}_{+}$ is the time-dependent particle distribution in phase space. The system’s evolution is driven by the total electric field, which is a sum of the externally applied control field $H(x)$ and the self-generated electric field $E_{f}(t,x)$. This self-generated field is determined self-consistently: the charge density $\rho_{f}(t,x)=\int f(t,x,v)\,\mathrm{d}v$ determines the electric potential $V_{f}(t,x)$ via the Poisson equation, and $E_{f}$ is determined by its gradient. The objective function $\mathcal{J}$ is a functional designed to quantify the degree of instability in the plasma distribution $f$.
This PDE-constrained optimization framework is a cornerstone of modern control theory and has been employed in numerous studies of plasma stability, both theoretical [18, 26, 27, 35, 34] and numerical [19, 1, 31, 4, 6, 7, 49]. Regardless of the specific physical system, the underlying optimization problem is invariably challenging. The nonlinear nature of the governing PDEs typically leads to highly non-convex objective landscapes, populated with numerous local minima. As such, the careful design of the objective function $\mathcal{J}$ is critical. Different choices for $\mathcal{J}$ emphasize different features of instability [29, 39]; for instance, one metric might penalize the kinetic energy of unstable wave modes, while another might measure the deviation of the entire distribution in phase space from its equilibrium state. This choice directly shapes the optimization landscape and determines which properties of the solution are promoted. This leads to the central question that this manuscript seeks to answer:
@@
What is a good metric for suppressing instabilities in the Vlasov-Poisson system?
In the context of fusion energy, the ultimate goal is to maintain the plasma’s energy in kinetic form, which implies minimizing the energy stored in the self-generated electric field [25, 37]. A common approach in the literature is to target this goal indirectly by minimizing the $L^{2}$-misfit between the final state of the distribution, $f_{T}:=f(T,\cdot,\cdot)$, and the desired equilibrium, $f_{\text{eq}}$ [35, 36]. The rationale behind this approach lies in the fact that if $f_{T}$ perfectly matches the spatially uniform equilibrium $f_{\text{eq}}$, the self-generated electric energy must be zero. However, as demonstrated numerically in [19], this $L^{2}$ objective function is highly non-convex and oscillatory, creating a landscape with a myriad of local minima that can trap optimization algorithms.
This motivates our search for alternative objective functions, which we explore mainly along two axes: the type of metric and its temporal structure. In the first axis, we compare two types of metrics:
  • To measure the deviation from equilibrium, we study using the Kullback-Leibler (KL) divergence as a metric. As $f$ is a probability distribution, the KL-divergence is a more natural information-theoretic measure of disparity than the $L^{2}$ norm.
  • To directly target the underlying physics, we propose minimizing the self-generated electric energy, a macroscopic quantity that is computationally efficient to evaluate and manipulate.
In the second axis, we investigate a crucial structural modification: we compare objective functions evaluated only at the final time $T$ against functions that are integrated over the entire time interval. This allows to test whether incorporating the full dynamics of the system can regularize the problem and appease the non-convexity of the optimization landscape.
However, our numerical landscape analysis reveals that even time-integrated objective functions, while possessing more favorable local properties, remain highly non-convex far from the global minimum. This implies that the success of any gradient-based optimization is critically dependent on the quality of the initial guess for the control field $H$.
We address this issue by analytically building a high-quality initial guess in which we leverage the rich analytical theory of the Vlasov-Poisson system [18]: we consider the linearized regime of the system around the equilibrium state and perform a stability analysis based on the dispersion relation. By identifying the fastest-growing unstable modes (via Fourier and Laplace analysis), we analytically derive a control field specifically tailored to suppress them. While this linear analysis is valid only for short times before non-linear effects dominate, we show numerically that the resulting control provides an initial guess that lies within the basin of attraction of the global minimum for the full non-linear problem (1.1).
In summary, this paper presents an integrated analytical-numerical methodology for suppresing instabilities in the Vlasov-Poisson system. Our numerical simulations, which systematically explore different objective functions, optimization solvers, and initialization strategies, lead to three principal findings:
  • First, we demonstrate that objective functions that integrate information over time yield significantly smoother and more convex-like optimization landscapes around the global minimum compared to those evaluated only at a terminal time.
  • Second, we find that macroscopic metrics such as the electric energy, while physically intuitive, can create spurious, unphysical local minima that trap optimizers, particularly adaptive-steps solvers such as line-search-based methods [41].
  • Third, we show that linear dispersion relation analysis provides a reliable and effective method for constructing an initial guess for the control, capable of placing the optimization within the global minimum’s basin of attraction.
This combined approach allows for the efficient and robust discovery of external fields that effectively suppress the instabilites.
1.1 Related work
Due to its significance, plasma control has attracted considerable research interest, leading to the development of a wide range of control strategies. As one might expect, these strategies are fundamentally shaped by the mathematical models used to describe plasma behavior. These models span a spectrum from macroscopic fluid descriptions—such as magnetohydrodynamics (MHD)[6, 7, 49]—to microscopic kinetic models like the Vlasov–Poisson system. Owing to the high mathematical and computational complexity of kinetic models, many approaches rely on simplifications such as dimensionality reduction[19] or linearization [39, 28, 18] to make the resulting control problems more tractable. These approximations allow the use of optimal control techniques within available computational resources.
Once a model is chosen, another key consideration is the nature of the control strategy itself. Broadly speaking, plasma control can be categorized into static and dynamic control. In static control—such as the approach pursued in this work—the system starts near equilibrium, with only a small perturbation that the control aims to dampen by a prescribed target time [19, 35, 36]. By contrast, dynamic control involves real-time feedback mechanisms that respond continuously to the evolving state of the plasma. Theoretical developments in this direction can be found in [18], while recent experimental implementations have emerged under the optimal control paradigm [1, 2, 7] and the reinforcement learning (RL) paradigm [16, 46]. These RL-based methods recast the control problem in a data-driven framework, training neural networks to predict instabilities [47] and develop reactive strategies. While RL offers the promise of real-time, adaptive control of nonlinear plasma dynamics, it also faces limitations—particularly the mismatch between actuation and measurement timescales (on the order of 25 ms [38]) and the much faster evolution of certain instabilities, which may occur on sub-millisecond or even nanosecond timescales.
Regardless of the control type, implementation in experimental or computational settings typically involves actuating either a magnetic field or an electric field. Magnetic field control is the dominant paradigm in magnetic confinement fusion experiments, particularly for managing MHD-scale behavior in tokamaks [6]. It has also been extended to kinetic models, where theoretical and numerical studies have investigated magnetic control within the Vlasov–Poisson framework to confine plasma away from reactor walls and prevent material damage [1, 2, 35, 36]. In contrast, electric field control—although less common in large-scale experimental devices—has proven powerful in theoretical and computational studies, particularly for kinetic models. Applications include beam shaping in particle accelerators [18] and, most relevant to this work, the suppression of kinetic instabilities via PDE-constrained optimization [19]. Moreover, in one-dimensional configurations, electric fields constitute the only physically viable external control mechanism.
Finally, we should stress that different mathematical definitions of “instability” (objective functions) show drastically different landscapes for a given shared physics problem is not a new phenomenon, and is especially pronounced for wave-type problems, see [20, 50, 13]. In the context of plasma control, though the result is expected, we have not found literature that thoroughly studied this phenomenon.
1.2 Organization
@@
The paper is organized as follows. In Section 2 we properly define the PDE-constrained optimization problem that we aim to solve, along the different possible objective functions that we will test. We also present the forward solver that we will use to solve the forward problem and the two canonical examples that we will present in this work. In Section 3 we briefly present the dispersion relation and a linear stability analysis which provides important information on the frequency and scaling that our control $H$ should have. In Section 4 we thoroughly study the different landscapes for different objectives functions. In Section 5 we make substantial numerical simulations with different objectives, methods, and initializations of the parameters.
@@ 2 Preliminaries
In practice, plasma for fusion applications is typically confined within toroidal or axisymmetric domains, such as those found in tokamaks or stellarators, which can be approximated to a circular domain. From a mathematical perspective, this corresponds to treating the $y,z$-domain as constant and expressing the 3D problem as a pseudo-1D problem with periodic boundary conditions, where $(x,v)\in[0,L_{x}]\times[-L_{v},L_{v}]$ with periodic boundary conditions in space and homogeneous Dirichlet boundary conditions in the velocity domain. Under this framework, the Vlasov-Poisson system can be simplified, resulting in the following formulation:
$\left\{\begin{array}[]{lc}\partial_{t}f+v\partial_{x}f-(E_{f}+H)\partial_{v}f=0\,,\\ E_{f}=\partial_{x}V_{f}\,,\\ \partial_{xx}V_{f}=1-\rho_{f}=1-\int f\,\mathrm{d}v\,.\end{array}\right$. (2.1)
@@
It is well known that equilibrium states which do not satisfy the Penrose condition [43] are classified as unstable. In such cases, even a small perturbation to the equilibrium state can lead to exponential growth of the self-generated electric field $E_{f}$. In this work, we study two canonical examples of such unstable equilibria that violate the Penrose condition [12, 40]: the Two Stream equilibrium (Section 2.3.1) and the Bump-on-Tail equilibrium (Section 2.3.2).
@@ 2.1 PDE-constrained optimization problem
$\min_{H}$ $\quad\mathcal{J}(f[H])$ (2.3)
@@
where our objective function will be either of the following four [^1] Another straightforward definition of objective function is to use the $L^{2}$ difference between the equilibrium and the solution $\mathcal{J}(f[H])=\frac{1}{2}\|f_{T}[H]-f_{\text{eq}}\|^{2}_{L^{2}(x,v)}=\frac{1}{2}\int_{-L_{v}}^{L_{v}}\int_{0}^{L_{x}}(f[H](T,x,v)-f_{\text{eq}}(v))^{2}\,\mathrm{d}x\,\mathrm{d}v\,$,. Numerically we find this objective function presents very similar landscape as the KL divergence, as will be discussed in Appendix A.
@@
$\mathcal{J}(f[H])=\mathcal{E}_{f[H]}(T)=\int_{0}^{L_{x}}[E_{f[H]}(T,x)]^{2}\,\mathrm{d}x\,$, ($E$)
$\mathcal{J}(f[H])=\text{KL}(f_{T}[H]||f_{\text{eq}})=\int_{0}^{T}\int_{-L_{v}}^{L_{v}}\int_{0}^{L_{x}}f[H](t,x,v)\log\left(\frac{f[H](t,x,v)}{f_{\text{eq}}(v)}\right)\,\mathrm{d}x\,\mathrm{d}v\mathrm{d}t\,$, (KLT)
$\mathcal{J}(f[H])=\int_{0}^{T}\mathcal{E}_{f[H]}(t)\mathrm{d}t=\int_{0}^{T}\int_{0}^{L_{x}}[E_{f[H]}(t,x)]^{2}\,\mathrm{d}x\,\mathrm{dt}\,$, (EET)
where $E_{f[H]}$ and $\mathcal{E}_{f[H]}$ are as defined in (2.1) and (2.2) respectively. We note that these four objective functions can be grouped into two categories: one based on electric energy and the other based on KL divergence. The former is a natural, physics-inspired choice, while the latter is a valid approach that has been considered in a recent work on control for the Vlasov-Poisson system [8].